OpenStudy (anonymous):

f(x) = x + 9, find the value of x if f(x) = 15. 15 24 6 9

OpenStudy (midhun.madhu1987):

x+9=15 x=?

OpenStudy (anonymous):

If f(x) = x + 9 and f(x) = 15, we can write this as 15 = x + 9 and hence x = 15 - 9 = 6